How can you prepare 0.25 mole of potassium nitrate from a stock with a concentration of 2mole . Using dilution ratio
1
Expert's answer
2020-01-24T05:00:02-0500

If there is 1 liter of 2M solution, then n(KNO3) = C1(KNO3) * V1(KNO3) = 2M * 1L = 2 mol.

V2(KNO3) = n(KNO3)/C(KNO3) = 2 mol/0.25M = 8L.

To prepare a 0.25M solution, you need to dilute the 2M solution 8 times.